TheArchitect:

Pt 8 find the slope of the line that passes through the pair of points. Write your answer as an improper fraction, if necessary. (−19, 45), (28, −32)

Aliciaa:

Answer: Slope m=−77/47 As a decimal: m = -1.6383

TheArchitect:

@aliciaa wrote:
Answer: Slope m=−77/47 As a decimal: m = -1.6383
Thank you Alicia 💟